OpenStudy (clamin):

HELP!! (x^3-2x^2-5x+6)/(x+2)

OpenStudy (anonymous):

x=1 x=3

OpenStudy (clamin):

solutions

OpenStudy (anonymous):

you can use factor theorem. f(x) = x^3+2x^2-5x-6 f(-1) = -1+2+5-6 = 0 x+1 is a factor use synthetic division, x^3+2x^2-5x-6 = (x+1)(x^2+x-6) = (x+1)(x-2)(x+3) = 0 therefore, the solution is x = -1, 2, -3

OpenStudy (clamin):

polynomial

Nnesha (nnesha):

yes